Find f ◦ g and g ◦ f, where f(x) = x
2 + 1 and g(x) = x + 2, are functions from R to
R
1
Expert's answer
2020-09-06T17:43:51-0400
"f(g(x))=f(x+2)=(x+2)^2+1"

"=x^2+4x+4+1"

"=x^2+4x+5"

Answer: x^2+4x+5



"g(f(x)) =g(x^2+1)=(x^2+1)+2"

"=x^2+3"

Answer: x^2+3